OpenStudy (anonymous):

find the first derivative sqrt[x^2-sqrt(2x-3)]

OpenStudy (shamil98):

chain rule

OpenStudy (shamil98):

try it

OpenStudy (anonymous):

is this right? sqrt[x^2-(1/2(2x-3)^-1/2 * 2)]?

geerky42 (geerky42):

We can let \(f(x) = \sqrt{x}\) and \(g(x) = x^2-\sqrt{2x-3}\), so we have \(f(g(x))\) to apply chain rule

geerky42 (geerky42):

\(\dfrac{d}{dx}f(~g(x)~) = f'(~g(x)~)~g'(x)\)

OpenStudy (anonymous):

or this 1/2*sqrt[x^2-sqrt(2x-3)]* d/dx (x2-sqrt 2x-3)

geerky42 (geerky42):

yes I think... hard to read O_o If you mean \(\Large \dfrac{1}{2\sqrt{x^2-\sqrt{2x-3}}}\cdot\dfrac{d}{dx}\left(x^2-\sqrt{2x-3}\right)\), then you are right

geerky42 (geerky42):

so finish derivative and put it in numerator, and you are done. i think you can simplify, but i'm sure teacher don't want you to so that

geerky42 (geerky42):

is that clear? @wakakaka

OpenStudy (anonymous):

yeah thanks a lot

geerky42 (geerky42):

welcome